### v2

# Non-Markovian Dynamics in Fiber Delay-Line Buffers

Abstract
We study the non-Markovian effect on a two-photon polarization entangled state in which one photon from the pair is stored in a fiber delay-line buffer. We propose a model of a photonic qubit coupled to fiber birefringence and a fiber reservoir representing the environment. We analytically derive a non-Markovian probability function for the buffered photon and its paired photon. To verify the probability function, we perform the full quantum state tomography of the photon pairs. The probability function fits well with the experimental data and the physical values. Our result indicates that our quantum system is operating slightly above the threshold for a non-Markovian transition. We observe a unique polarization dynamic of the buffered photon. We further exploit the measures of quantum mutual information for studying the quantumness of the photon pairs. We find that Werner’s well-known separability criterion occurs at the buffer time of about 0.9 ms. Our result implies that the quantum discord can surpass Werner’s criterion, and hence the quantum bi-partite correlation can exist for a buffer time greater than 0.9 ms.
pacs:
03.67.Hk, 42.50.Lc, 03.67.Mn, 42.50.Ar
I Introduction
Photons that traverse in optical fiber with little attenuation are ideal carriers for quantum networking. The photonic qubits with entanglement in polarization, time-bin, or other degrees of freedom are the resources for establishing quantum communication and distributed quantum computing. Quantum buffers made of optical fiber and free-space optics can provide the entire telecom C-band optical bandwidth and storage time up to $\mu\rm{sec}$ without significant deterioration of storage efficiency kflee24 ; pittman02 ; pittman23 ; XLi_OL_2005 ; Kwiat_SPIE_2021 ; Kwait_optica_2017 ; Wang_2022 ; Takesure_ncomm_2013 ; Park_OE_2019 ; Clemmen_OL_2018 ; Mendoza_Opa_2016 . Short buffer times (few nanoseconds) can provide active temporal source multiplexing for high-efficiency single photon generation Park_OE_2019 ; Kaneda_2015 ; Xiong_NC2016 . Long storage times play an important role in optical networks for packet synchronization, label processing, and contention management. Quantum buffers based on fiber delay lines are traveling buffersZhong_2001 ; Kaneda_2015 ; Xiong_NC2016 and fiber loop buffers Weber_1996 ; kflee24 . The traveling buffers, which can provide qubit storage time up to $\mu\rm{sec}$, are relatively unexplored. In general, there is a strong coupling between qubits, fiber birefringence, and a fiber reservoir, so we need to develop a non-Markovian model to study the decoherence dynamics of qubits in a fiber-based quantum buffer.
Recent studies Giovannetti_PRL_22 ; Giovannetti_JPA_05 ; Helsen_2023 indicate that Markovian and non-Markovian processes can be controlled for improving the quantum capacity in a high-loss communication channel and modeling quantum networks. Interestingly, one can protect quantum entanglement by using the memory effect of the classical noises in the non-Markovian regimeWenjie_2023 . The fiber-based quantum buffer is an open quantum system because the polarization degree of freedom of photon is coupled with the frequency degree of freedom of the fiber birefringence which is coupled with the fiber reservoir. The density matrix of a qubit can evolve into Markovian (non-Markovian) dynamics depending on the weak (strong) coupling with the medium, respectively. The medium here is a single-mode optical fiber birefringence and a fiber reservoir. In a weak coupling process, the medium induces memoryless dynamics on the qubit so that its density matrix $\rm{\rho(t)}$ at time t is determined by $\rm{\rho(t-dt)}$ with the infinitesimal $dt$. In other words, the information of a qubit should not be memory-related with the previous qubit. The Markovian dynamic of a qubit is ideal for long-distance communication networks. The non-Markovian process, in contrast, keeps the memory of the coherence dynamics of the qubit such that its density matrix $\rm{\rho(t)}$ at time t is determined by its past $\rm{\rho(t^{\prime})}$ where $t^{\prime}$ is the start of the strong coupling process. It is essential to minimize the non-Markovian dynamic of a qubit stored in a fiber-based buffer by exploring its coupling with fiber birefringence and fiber reservoir. A comprehensive discussion of concepts, measures, and witness of non-Markovian effect for an open quantum system can be found in refPlenio_2014 ; Piilo_16 . A trace distance of two quantum states Piilo_09 ; Piilo_10 is usually used to detect the crossover between the Markovian and non-Markovian processes such as the coupling between the photon polarization and its optical frequency degree of freedom in a simple titled Fabry-Perot cavity platform Guo_nphy_2011 . The non-Markovian effect can revive entanglement after observation of entanglement sudden death Haase_2018 ; Wang_2018 .
An exactly solvable master equation of an open quantum system, such as a two-level atom interacting resonantly in a reservoir formed by the Lorentzian quantized modes of a high Q cavity, is often used as an example for illustrating the non-Markovian dynamics of a qubit Breuer_oxford_2002 ; Petruccione_PRA_2006 ; Bellomo_PRL_2007 ; Chen_PRA_14 ; Wang_scr_17 . Furthermore, the non-Markovian dynamics have been studied by using quantum discord Wang_PRA_2010 ; Fanchini_PRA_2010 , which is another concept of quantum correlation introduced by quantum mutual information theory. Some believe that quantum discord is more suitable for quantifying the quantum correlation of a bipartite system in non-Markovian environments. We recently used it to explore the quantum coherence of a two-qubit entangled state generated in green fluorescent protein Kflee_nc_2017 ; Shi_sr_2016 at room temperature. One can experimentally obtain quantum discord by calculating the total correlation and classical correlation based on the density matrix elements of a quantum state.
In this paper, one photon from the polarization entangled photon pair is stored in the buffer. We use a model to study the non-Markovianity of a photonic qubit coupled to the fiber birefringence of a single-mode optical fiber in a fiber reservoir. We analytically derive the probability function for the buffered polarization qubit and its paired photon. Quantum state tomography (QST) is then performed on the entangled photon pairs. We obtain the probability function from the reconstructed density matrix. Our model agrees with the experimental data. The buffered photon exhibits the non-Markovianity because of the polarization modal dispersion (PMD) and the loss in the fiber reservoir. In the experiment, we also prove the existence of the Markovian process in the scenario where we replace the buffer with a free-space attenuator.
II Results
Figure 1: A simple model consists of a photonic qubit, fiber birefringence in optical fiber, and a fiber reservoir representing the environment decoherence. The $\kappa$ is the coupling constant between the qubit and fiber birefringence ($\Delta n$). The $\kappa(\rm{sec}^{-1})$ is inversely proportional to PMD ($\sim\frac{1}{\sqrt{\langle\tau^{2}\rangle}}$), where $\langle\tau^{2}\rangle$ is differential group delay (DGD). The fiber birefringence is then coupled with a fiber reservoir with the coupling constant $\gamma_{\circ}(\rm{sec}^{-1})$ which is responsible for the loss of qubit information such as decoherence.
Here, we consider a qubit-fiber-reservoir model consisting of a photonic qubit, a single-mode optical fiber, and a fiber reservoir representing the environment as shown in Fig.1. (see Methods for the atom-cavity-reservoir model). When a photon propagates in a long optical fiber, its polarization state and transmission rate will be degraded by the fiber birefringence and a fiber reservoir. The fiber reservoir is the medium where a polarization qubit interacts/exchanges energy with the environment. The polarization of the photon with certain optical frequency bandwidth is coupled with the fiber birefringence which is coupled with the fiber reservoir. In other words, the PMD effect is coupled to the environment via the fiber reservoir. The photon polarization state will experience the dephasing caused by the PMD and the absorption/scattering process in the fiber. The absorption and re-emission of a photon state in the fiber will decohere a quantum state. A photonic qubit with polarization state H or V is coupled with the fiber birefringence by a coupling constant $\kappa\propto\frac{1}{\rm{PMD}}$. A photonic qubit in a single optical frequency mode will couple weakly with the fiber birefringence or less dispersion (large $\kappa$), and hence experiencing less depolarization effect. If we increase the frequency bandwidth of the photonic qubit in a steep dispersion region, the $\kappa$ becomes small. The qubit will start to experience the PDM effect, which is given by the square root of time variance $\langle\tau^{2}\rangle$ of differential group delay (DGD) Kogelnik_PMD_2002 ; Wai_jlt_1996 ; Brodsky_jlt_2006 or $\sqrt{\langle\tau^{2}\rangle}$. Therefore, the strength of the PMD effect is inversely proportional to the $\kappa$. In other words, one can enhance the PMD effect on the photonic qubit by increasing their differential group delays such as inserting a few polarization-maintaining fiber jumpers in their channels. The same photon with optical frequency $\omega$ can be spontaneously absorbed and scattered at a rate of $\gamma_{\circ}$ due to the coupling between the fiber birefringence and the reservoir. The $\frac{\gamma_{\circ}}{c}\propto\mu$ is the loss parameter per unit length of the photon due to photon absorption/scattering in the fiber reservoir. (see Methods for the probability function of the photon state and how the relative strength of the $\gamma_{\circ}$ and $\kappa$ can determine the photon state undergoing Markovian and non-Markovian processes). Two factors contribute to the loss parameter. The first is the propagation loss of the photonic qubit in the fiber. The second is the loss of information on the photonic qubit which is its polarization property. The environment depolarization process can degrade the polarization extinction ratio of the qubit. The propagation loss of the photonic qubit can cause an increase in the accidental coincidence and reduce the fidelity of the quantum state. In this paper, we subtract the accidental coincidence of the signal and idler so that we can isolate the contribution of the propagation loss to the reconstructed density matrix and study the polarization property of the buffered qubit which is the idler photon.
Figure 2: a) H(V) analog to the excited state (ground state) of a two-level atomic system. The geometry picture for the change of polarization state (i) $\rm{H}\rightarrow\rm{H}^{f}$ and (ii) $\rm{V}\rightarrow\rm{V}^{f}$. Small arrows along the $\rm{H}^{f}$ and $\rm{V}^{f}$ show the contribution from the H and V. The phase fluctuation $\Delta\phi$ due to PMD degrades the polarization extinction ratio of the idler. b) Experiment setup including the buffer made of two low-loss switches with a short path and a long path. We use different fiber lengths in the long path to change the buffer time.
In our two-qubit system, the photon pair is initially prepared in the entangled state $|\psi_{\circ}\rangle=\frac{1}{\sqrt{2}}[|H_{s}H_{i}\rangle+|V_{s}V_{i}\rangle]$. One photon from the pair is sent to a buffer with the fiber length of L. Let’s say the photon of interest is the idler which is experiencing the PMD and the loss in the buffer. The initial state of the idler (signal) is $|H_{i}\rangle+|V_{i}\rangle$ ($|H_{s}\rangle+|V_{s}\rangle$), respectively. Due to depolarizing caused by the PMD and fiber reservoir, the final components $H^{f}_{i}$ and $V^{f}_{i}$ can be written in terms of $H_{i}$ and $V_{i}$ as
$H^{f}_{i}=H_{i}\rm{cos(\Delta\phi_{h})}+V_{i}\rm{sin(\Delta\phi_{h})}$
$V^{f}_{i}=V_{i}\rm{cos(\Delta\phi_{v})}-H_{i}\rm{sin(\Delta\phi_{v})}$ (1)
where $\Delta\phi_{h(v)}$ is the amount of phase change caused by the PMD effect. The phase change $\Delta\phi_{h(v)}=\Delta\omega\,\sqrt{\langle\tau^{2}\rangle}$ is related to the square root of time variance $\langle\tau^{2}\rangle$ of differential group delay (DGD) Kogelnik_PMD_2002 ; Wai_jlt_1996 ; Brodsky_jlt_2006 or $\sqrt{\langle\tau^{2}\rangle}\propto L^{1/2}$, where L is the length of the fiber buffer. The $h$ ($v$) sign denotes the phase change associated with the polarization state $H_{i}$ ($V_{i}$). When the $\Delta\phi_{h}$ and $\Delta\phi_{v}$ are zero, the $H^{f}_{i}\rightarrow H_{i}$ and $V^{f}_{i}\rightarrow V_{i}$. The polarization states of the idler do not change because of the absence of the PMD effect. When the $\Delta\phi_{h}$ and $\Delta\phi_{v}$ are +90(90)9090+90(-90)+ 90 ( - 90 ), $H^{f}_{i}\rightarrow V_{i}(-V_{i})$ and $V^{f}_{i}\rightarrow-H_{i}(+H_{i})$, respectively. The polarization states of the idler are completely flipped. The geometrical meaning for these notations is shown in Fig.2(a). The phase fluctuation of $\rm{cos(\Delta\phi)}$ and $\rm{sin(\Delta\phi)}$ degrade the polarization extinction ratio of the idler. One can interpret that the final idler polarization state $H^{f}_{i}$ is related to its initial states $H_{i}$ and $V_{i}$ in the past because of the non-Markovian effect caused by the PMD. We assume that the signal is propagating in a short fiber length before it is detected. We then obtain the probability function $\mathcal{P}_{f}=|\langle\psi_{\circ}|\psi^{f}\rangle|^{2}$ as,
$\mathcal{P}_{f}=|\mathcal{A}[\rm{cos(\Delta\phi_{h})}+\rm{sin(\Delta\phi_{h})}-\rm{sin(\Delta\phi_{v})}+\rm{cos(\Delta\phi_{v})}]|^{2}$ (2)
where $\mathcal{A}\mathcal{A}^{*}=e^{-2\mu L}$ is introduced due to the loss parameter in the fiber reservoir on the polarization properties experienced by the surviving idler photons. In our experiment, we diminish the contribution of the propagation loss to the probability function by subtracting the accidental coincidence, so that we can measure the probability function which is degraded by the $\mathcal{A}\mathcal{A}^{*}=e^{-2\mu L}$ due to the fiber reservoir.
We explore non-Markovian dynamics of the entangled state by simplifying Eq.(2) into two cases; the magnitude of $\Delta\phi_{h}=\Delta\phi_{v}$ and $\Delta\phi_{h}\neq\Delta\phi_{v}$. For the case $\Delta\phi_{h}=\Delta\phi_{v}=\Delta\phi$, we have the first scenario where the idler polarization states $H_{i}$ and $V_{i}$ rotate in opposite directions i.e we substitute the terms $+\Delta\phi_{h}$ and $-\Delta\phi_{v}$ or $-\Delta\phi_{h}$ and $+\Delta\phi_{v}$ in Eq.(2). The $+$ (--) value of the $\Delta\phi_{h(v)}$ corresponds to the rotation of the polarization state in the clockwise (anti-clockwise) direction. The probability function $\mathcal{P}_{f}$ is simplified as
$\mathcal{P}_{a}\propto e^{-2\mu L}[\rm{cos(\Delta\phi_{1})}\pm\rm{sin(\Delta\phi_{1})}]^{2}$ (3)
where $\Delta\phi_{1}=\Delta\omega\,\sqrt{\langle\tau_{1}^{2}\rangle}$. Eq.(3) has a similar form as Eq.(8) in Methods for a two-level atom interacting in a high Q-cavity with a memoryless reservoir. This implies that the idler is experiencing the non-Markovian process due to the coupling between the PMD effect and the loss in the fiber reservoir. Let’s consider the second scenario where the idler polarization states $H^{f}_{i}$ and $V^{f}_{i}$ rotate in the same directions i.e, we substitute the terms $+\Delta\phi_{h}$ and $+\Delta\phi_{v}$ or $-\Delta\phi_{h}$ and $-\Delta\phi_{v}$ in Eq.(2). Then the probability function $\mathcal{P}_{f}$ becomes
$\mathcal{P}_{sy}\propto e^{-2\mu L}[\rm{cos(\Delta\phi_{2})}]^{2}$ (4)
where $\Delta\phi_{2}=\Delta\omega\,\sqrt{\langle\tau_{2}^{2}\rangle}$. The Eq.(4) has the same form as Eq.(9) in Methods for large $n$. For the case $\Delta\phi_{h}\neq\Delta\phi_{v}$, the complete probability function is shown in detail in the Method.
Later, we will verify the probability functions of Eq.(3) and Eq.(4) and also Eq.(8) and Eq.(9) in the Methods by launching one photon from the entangled photon pair in a quantum buffer made of different fiber lengths as shown in Fig.2(b). We use a counter-propagating scheme (CPS) Kflee_ol_2006 ; Sua_ol_2014 to generate two-photon polarization entangled state $|\psi_{\circ}\rangle$ through a four-wave mixing process in a 300 m of dispersion-shifted fiber. The detailed experimental setup is shown in Fig.1S in the Supplementary. The signal and idler photons are separated by the DWDMs (dense wavelength-division multiplexing). The signal photon is sent to the polarization analyzer $\rm{PA_{s}}$. The idler is sent to the buffer which consists of two low-loss switches (insertion loss $\leq$ 0.01dB and a single-mode optical fiber as shown in Fig.2(b). We can select the idler to pass a long or short fiber length. When the idler is selected to pass the long fiber, then the idler is buffered with the duration of $\frac{L}{c/n_{r}}$, where $L$ is the length of the fiber and $n_{r}$ is the refractive index of the fiber. The idler is then directed to the polarization analyzer $\rm{PA_{i}}$. The first pair of the quarter-wave plate (QWP) and half-wave plate (HWP) is used to compensate for the birefringence of the fiber. The second pair of QWP and HWP is used to implement quantum state tomography. The timing of detecting the buffered photon can cause a non-Markovian effect on the photon stateHelsen_2023 ; Humprey_Nature_18 . In this buffer configuration, the qubit storing process is initialized by two units of 1$\times$2 switch. This switching process can impose loss and decoherence such as the amplitude/phase damping process. We use two units of 1$\times$2 switch instead of one unit of 2$\times$2 switch because the insertion loss of 1$\times$2 switch (2$\times$2 switch) is about $<$ 0.01 dB (1.2 dB), respectively. To avoid decoherence due to timing issues on the photon detection, we operate the switch with low speed i.e. 1.0Hz. The slow switching rate allows us to take data when the switches are both turned on. So, we can exclude the data appears in the rising and falling edges of the switches. In other words, we diminish the non-Markovian effect which can be caused by the switching process kflee24 . We use an interface program to turn on and off the two switches synchronously at the rate of 1.0 Hz. The data is taken on the duration when both switches are completely turned on/off. We then perform quantum state tomography (QST) by using two single-photon detectors (NuCrypt CPDS-4) operated at 50 MHz.
We characterize the prepared state $|\psi_{\circ}\rangle=(|H_{s}H_{i}\rangle+|V_{s}V_{i}\rangle)/\sqrt{2}$ by using the standard method of QST James_pra_2001 ; White_pra_2001 ; Munro_pra_2001 . There are a total of 16 settings for the HWPs and QWPs in the $\rm{PA_{s}}$ and $\rm{PA_{i}}$. For each setting, we measure the coincidence counts (CC) and accidentals (AC) with an integration time of 2 seconds or 100M sampling gates when during the switches are turned on. Using the maximum likelihood estimation, we reconstruct the $4\times 4$ density matrix of the final state $\rho^{f}$ in the HV basis. We construct the density matrix $\rho^{f}$ based on the true coincidence counts i.e. the accidental coincidence is subtracted. In other words, the Raman photon and loss-induced accidental do not contribute to this study. We repeat the state tomography for different fiber lengths from 0.0 m up to 25 km. The 0.0 m fiber length corresponds to the short path in the buffer.
We use the quantum discord to characterize the non-Markovian dynamics of the buffered photon. It is of both fundamental and practical interest to differentiate classical and quantum correlations within the total correlation for the buffered photon experiencing a non-Markovian environment. Two scenarios can contribute to the classical information; (1) environment projects or makes a measurement on the idler qubit state and creates the classical state, (2) environment perturbation can dechore the information of the idler qubit state such as degradation of the polarization properties. Scenario (1) is the classical correlation which will appear at the coincidence and accidental coincidence. We can avoid this classical correlation by subtracting the accidental. Scenario (2) can create a mixture of product states. We study the quantum discord under scenario (2) which can allow us to have a buffer time $>$ 0.9 ms. For this purpose, we can accurately treat the final state as the Werner state. The Werner state is the mixture of the maximally entangled state with the probability $P$ and the maximally mixed state as given by Werner_pra_1989 ; Munro_pra_2001 ,
$\mathcal{W}=P|\psi_{\circ}\rangle\langle\psi_{\circ}|+\frac{1-P}{4}I.$ (5)
where the $I$ is the identity matrix. We calculate the average of the probability $P=\sum^{n=6}_{i}p^{\prime}_{i}/n$, where $p^{\prime}_{1(2)}=4\rho^{\prime}_{11(44)}-1$, $p^{\prime}_{3(4)}=2\rho^{\prime}_{14(41)}$, and $p^{\prime}_{5(6)}=1-4\rho^{\prime}_{22(33)}$. The density matrix elements $\rho^{\prime}_{11}$, $\rho^{\prime}_{44}$, $\rho^{\prime}_{14}$, $\rho^{\prime}_{41}$, $\rho^{\prime}_{22}$, and $\rho^{\prime}_{33}$ are obtained from the final state. The measured average probability $P$ is the non-Markovian probability function p(t). ( See Method for Amplitude damping on one photon from two-photon Werner state).
We plot the average probability $P$ of the Werner state as a function of buffer time $t$ as shown (red solid circle) in Fig.3. We fit the experimental data with the summation of the probability functions $\mathcal{P}_{a}$ of Eq.(3) and $\mathcal{P}_{sy}$ of Eq.(4). With the substitution of $L=(c/n_{r})t$ and $\sqrt{\langle\tau^{2}\rangle}=D_{p}\,\sqrt{L}$, the $\Delta\phi_{1,2}$ in the harmonic functions of $\mathcal{P}_{a}$ and $\mathcal{P}_{sy}$ is then given by a $\Delta\phi_{1,2}=\Delta\omega\,\sqrt{\langle\tau_{1,2}^{2}\rangle}=\Delta\omega\,D_{p1,p2}\,\sqrt{\frac{c\cdot t}{n_{r}}}$, respectively, where $D_{p1,p2}$ is the PMD parameter in unit of $\rm{picosecond}/\sqrt{km}$. Let’s denote $\mathcal{P}_{a,sy}=\mathcal{P}_{a}+\mathcal{P}_{sy}$. The good fit of $\mathcal{P}_{a,sy}$ exhibits a small modulation on the exponential function of $e^{-2\mu\frac{c}{n_{r}}t}$ as shown (red line) in in Fig.3. The modulation is caused by the harmonic function of $\sqrt{t}$ in $\mathcal{P}_{a,sy}$. This implies that the probability $P$ of the Werner state exhibits non-Markovian dynamic. We substitute the experiment value of $\Delta\omega=2\pi\times 200\rm{GHz}$ in $\mathcal{P}_{a,sy}$. We obtain $D_{p1}=0.0017\,ps$ and $D_{p2}=0.047\,ps$. Then, the total PMD parameter $D_{p}=D_{p1}+D_{p2}=0.05\,ps$ which is in agreement with the PMD link value of $\leq 0.1\frac{ps}{\sqrt{km}}$ (Corning SMF-28 Optical Fiber data sheet). We also obtain the loss parameter $\mu=6.0\times 10^{-6}\,m^{-1}$ from the fitting of the experimental data. We can rewrite the loss parameter $\mu$ as 0.006/km or 0.025 dB/km. This value is below 0.05 dB/km which is the induced attenuation caused by the environmental effects such as temperature and humidity.
We also fit the experimental data with the summation of the probability functions $p_{1}(t)$ of Eq.(8) and $p_{2}(t)$ of Eq.(9) in Methods. Let’s denote $p_{3}(t)=p_{1}(t)+p_{2}(t)$. The $p_{3}(t)$ exhibits stronger oscillation than $\mathcal{P}_{a,sy}$ within the error bar of the data as shown (blue line) in Fig.3. The equation of $p_{3}(t)$ indicates stronger non-Markovian dynamics than the $\mathcal{P}_{a,sy}$ because the harmonic function of $p_{3}(t)$ is a function of $t$. We obtain the parameters $\kappa_{1}=753\,{s}^{-1}$, $\kappa_{2}=3528\,{s}^{-1}$, and $\gamma_{\circ}=16292\,{s}^{-1}$ from the fitting of the experimental data. The total $\kappa=4281\,{s}^{-1}$. We have $4\times\kappa=17124\,{s}^{-1}$ greater than $\gamma_{\circ}$ satisfying the non-Markovian criterion i.e. $4\kappa\geq\gamma_{\circ}$. Our experimental data confirms the non-Markovian dynamic of the buffered qubit.
We further prove the existence of the Markovian process if the idler photon is not buffered. We replace the buffer in the idler channel with a free-space attenuator and adjust the attenuation $\alpha$ such that it equals the propagation loss (0.2 dB/km) $\times$ the fiber length. We then perform the quantum state tomography on the signal and idler. We obtain the average probability $P$ by using the same approach as discussed above. We plot the probability $P$ (blue data point) and also fit the data with any arbitrary exponential function as shown (blue dotted line) as shown in Fig.3. This indicates that the idler in free-space configuration is experiencing the Markovian dynamic without the fiber buffer i.e. without the coupling to the fiber birefringence and the fiber reservoir.
Figure 3: The measured probability $P$ of the Werner state as a function of buffer time for the buffer experiment (red solid circle) and the free-space attenuator (blue solid square). We convert the amount of the free-space attenuation ($\times$ 1/(0.2dB/km)) to buffer length and then to buffer time. The blue line (red solid line) is the fitting curve of the equation $p_{3}(t)$ (the function $\mathcal{P}_{a,sy}$), respectively. The blue dotted line is the fitting curve of an exponential function for the free-space attenuation.
Since the probability functions $\mathcal{P}_{a,sy}$ and $p_{3}(t)$ fit well with the measured probability $P$ of the Werner state, we can use the $\mathcal{P}_{a,sy}$ and $p_{3}(t)$ for quantum mutual information to quantify the quantum and classical correlation of the buffered photon and its paired photon. For a bipartite system, the total correlation, which is a measure of quantum mutual information, is given by $\mathcal{I}(\rho^{f})=\mathcal{C}(\rho^{f})+\mathcal{Q}(\rho^{f})$, where the $\mathcal{C}(\rho^{f})$ is classical correlation and the $\mathcal{Q}(\rho^{f})$ is quantum discord. For the Werner state, the quantum discord is given by,
$\mathcal{Q}(\rho^{f})=\mathcal{I}(\rho^{f})-\mathcal{C}(\rho^{f})$ (6)
where the total correlation and the classical correlation are given by Luo_pra_2008 ; Donnert_nmeth_2007 ,
$\mathcal{I}(\rho^{f})=\frac{3(1-P)}{4}\rm{log}_{2}(1-P)+\frac{(1+3P)}{4}\rm{log}_{2}(1+3P)$
$\mathcal{C}(\rho^{f})=\frac{(1-P)}{2}\rm{log}_{2}(1-P)+\frac{(1+P)}{2}\rm{log}_{2}(1+P)$ (7)
We also obtain the concurrence $\mathcal{C}_{n}(\rho^{f})=\rm{max}\{0,\frac{(3P-1)}{2}\}$ for the Werner state in the system.
Figure 4: The experimental data: quantum discord (blue triangle), classical (red triangle), and concurrence (orange circle). The quantum discord $\mathcal{Q}$ (blue line and blue dotted line), classical $\mathcal{C}$ (black line and black dotted line), and concurrence $\mathcal{C}_{n}$ (orange line and orange dotted line) of the Werner state as a function of buffer time by using the probability function $\mathcal{P}_{a,sy}$ and $p_{3}(t)$, respectively.
We plot the quantum discord $\mathcal{Q}(\rho^{f})$, classical correlation $\mathcal{C}(\rho^{f})$, and concurrence $\mathcal{C}_{n}$ as a function of the buffer time as shown in Fig.4. The experimental data is calculated based on the Eq.(6) where the probability $P$ is obtained from the measured density matrix of the Werner state. We also calculate the quantum discord, classical, and concurrence by using the $P=\mathcal{P}_{a,sy}$ (solid line) and the $p_{3}(t)$ (dotted line) as shown in Fig.4. See Fig.S2 in the Supplementary for the extrapolation plot of the $\mathcal{I}(\rho^{f})$, $\mathcal{C}(\rho^{f})$, $\mathcal{Q}(\rho^{f})$ and concurrence $\mathcal{C}_{n}$ as a function of the buffer time up to 1.5 ms. When $P>\frac{1}{3}$ which is the well-known Werner’s criterion for inseparable, the concurrent is not zero, so the Werner state is still entangled or inseparableZurek_prl_2001 . When the $P=\frac{1}{3}$, the concurrent is 0. This occurs at the buffer time = 0.9 ms (190 km) which is numerically obtained from the fitting curve of the function $\mathcal{P}_{a,sy}=\frac{1}{3}$ (see Fig.S2 and Fig.S3(a) in the Supplementary). As for the function $p_{3}(t)=\frac{1}{3}$, this falls at the buffer time = 0.4 ms (80 km) (see Fig.S2 and Fig.S3(b) in the Supplementary). The strength of the non-Markovian effect can determine the buffer time (fiber length) where the buffered photon can still preserve quantum entanglement. Since the function $\mathcal{P}_{a,sy}$ is a better fit, our data indicates that the fiber-based quantum buffer can preserve quantum entanglement with a buffer time of up to 0.9 ms.
Another interesting feature portrayed by the Werner state is the relationship between its concurrence, quantum discord, and classical correlation. For the function $\mathcal{P}_{a,sy}$ in the range $10.523<\mathcal{P}_{a,sy}<$ 1, the concurrence is larger than the quantum discord and classical correlation for the buffer time (length) from 0 s(0 m) to 0.7 ms(145 km). The $\mathcal{P}_{a,sy}=0.532$ occurs at the buffer time = 0.7 ms as shown in Fig.S2 and Fig.S3(a) in the Supplementary. Similarly, for the function $p_{3}(t)=0.532$, this falls at the buffer time = 0.3 ms (64 km) as shown in Fig.S2 and Fig.S3(b) in the Supplementary. The most important feature of the quantum mutual information is the quantum discord $\mathcal{Q}(\rho^{f})$ $>$ the concurrence and the classical correlation. This occurs at the buffer time larger than 0.7 ms(0.3 ms) for the function $\mathcal{P}_{a,sy}$ ($p_{3}(t)$), respectively.
III Discussion
To our knowledge, this measurement is the first demonstration based on a simple qubit-fiber birefringence-reservoir model to study the non-Markovian effect on a two-photon polarization entangled state in which one photon from the pair is stored in a buffer made of different fiber lengths. We compare our model with the traditional atom-cavity-reservoir model. The function $\mathcal{P}_{a,sy}$ fits the experimental data better than the $p_{3}(t)$. This is because the harmonic function of $\mathcal{P}_{a,sy}$ is a function of $\sqrt{t}$ due to the PMD effect in our model. While the harmonic function of $p_{3}(t)$ is a function of $t$ (see Methods for atom-cavity-reservoir model). We first fit our experimental data with the $\mathcal{P}_{a,sy}$ which is the summation of Eq.(3) and Eq.(4). We find that the contribution from Eq.(4) is more dominant than the Eq.(3). This implies that the H and V polarization components of the buffered photon rotate in the same direction. This finding is a new insight into the polarization properties of the buffered qubit that undergoes the non-Markovian dynamic. We then fit our experimental data with the $p_{3}(t)$ which is the summation of Eq.(8) and Eq.(9). The obtained $\kappa$ is slightly above the threshold $\frac{\gamma_{\circ}}{4}$ which is the condition for an open quantum system to become non-Markovian. Since our quantum system is operating close to the threshold $\kappa\sim\frac{\gamma_{\circ}}{4}$, the configuration of our experiment presents the possibility of naturally occurring crossover between the Markovian and non-Markovian processes.
We subtract the accidental coincidence in our data analysis so that the loss parameter $\mu$ is mainly contributed from the degradation of the polarization extinction ratio of the photonic qubit. One can regard the $\mu$ as the strength of the environment depolarization effect on the buffered photon. The obtained value of $\mu$ is below the induced attenuation value 0.05 dB/km caused by environmental factors such as temperature and humidity. Our finding indicates that the density matrix $\rho(t)$ of the entangled state at the output of the buffer is related to its early $\rho(t^{\prime})$ in the buffer, where $t>t^{\prime}$. This occurs when the environment depolarization effect of the fiber is more dominant than the PMD effect. We also demonstrate the Markovian scenario where the PMD effect did not exist by replacing the buffer with a free-space attenuator representing a memory-less environment (fiber birefringence and fiber reservoir).
We use the Werner state as the final state for the buffered idler photon and its paired signal photon. The Werner state is entangled (inseparable) when the $P>\frac{1}{3}$. The concurrence is zero at $P=\frac{1}{3}$ which is the Werner’s well-known separability criterion. The entanglement is related to the concurrence as $E_{n}(\rho_{AB})=H((1+\sqrt{1-\mathcal{C}_{n}^{2}})/2)$, where the $\mathcal{C}_{n}$ is the concurrence, and the function $H(x)=-x\log_{2}x-(1-x)\log_{2}(1-x)$ is the Shannon entropy. When the concurrence is zero, the entanglement is also zero at $P=\frac{1}{3}$. Based on this criterion, we can design the optimum fiber length of up to 190 km for the buffer in which the state is still inseparable or entangled. This buffer length corresponds to the buffer time of 0.9 ms.
Several measures of detecting the non-Markovian dynamics have been proposedLuo_pra_2012 ; Plenio_PRL_2010 ; Piilo_16 ; Piilo_09 ; Piilo_10 . The most recent method uses quantum process capabilities of non-completely-positive (CP) processes Chan_24 for identifying the non-Markovian dynamics. Our approach here of using quantum state tomography to find the probability function $p(t)$ leads us to the measures of quantum mutual information such as classical correlation and quantum discord. We then use the quantum discord to characterize the quantum correlation of the entangled state in our buffer system. Quantum discord is a better concept than entanglement to characterize the quantum correlation of an open quantum system Zurek_prl_2001 . The quantum discord is nonzero at $P=\frac{1}{3}$ where the concurrence is zero. (see Fig.S3 in the Supplementary). The non-zero quantum discord can surpass Werner’s criterion on the optimum buffer time. The quantum discord implies that the quantum bi-partite correlation still exists for buffer time greater than 0.9 ms.
The possibility for the buffered photon to experience the case $\Delta\phi_{h}\neq\Delta\phi_{v}$ is rare for the single-mode fiber which is used for communication networks. The reason is that the distributed fiber has a homogenous refractive index. The PMD effect of a light field in the single-mode fiber should not depend on the input polarization state of the light. However, the case $\Delta\phi_{h}\neq\Delta\phi_{v}$ may occur for the photon buffered in a loop of fiber more than one round trip. We are currently investigating the quantum buffer made of a fiber loop in a 2x2 Switch configurationkflee24 , where the photon is buffered by passing the same fiber loop multiple times.
Our analytical approach to derive the probability function can be easily extended to two qubits interacting separately with two independent non-Markovian environments such as the idler and the signal are buffered independently in two different buffers. It has been demonstrated that the non-Markovian density matrix evolution on one qubit state can mathematically proceed to the density matrix of a two-qubit entangled state Breuer_oxford_2002 ; Petruccione_PRA_2006 ; Bellomo_PRL_2007 ; Bellomo_PRA_2008 . The strength of the non-Markovian effect will be proportional to the square of Eq.(3) and Eq.(4), and hence will reduce the optimum buffer time. However, one can design the parameters $\kappa$ and $\gamma_{\circ}$ so that the crossover between Markovian and non-Markovian on the qubits can be controlled and utilized for modeling quantum networks Helsen_2023 ; Plenio_2014 and protecting quantum entanglementWenjie_2023 ; Franco_23 . We may improve the fiber-based quantum buffer by using the ultra-low loss single mode fiber such as SMF-28-ULL (Corning) and low loss Z-fiber from Sumitomo Electric, so that we can control the buffer to exhibit the Markovian effect, where its $\rho(t)$ is not related to $\rho(t^{\prime})$ in the buffer or the buffered qubit should not have memory effect on the previous/later qubit.
IV Methods
Atom-Cavity-Reservoir Model. Our analytical approach is similar to a qubit of a two-level atom interacting in a high Q-cavity, which is coupled to the reservoir with the Lorentzian spectral distribution Chen_PRA_14 ; Wang_scr_17 . The cavity-reservoir coupling gives rise to the effective spectral density in the form of, $(\gamma_{\circ}\Lambda^{2})/[(\omega_{\circ}-\omega)^{2}+\Lambda^{2}]$, where the parameter $\gamma_{\circ}$ is the coupling constant between the cavity and the reservoir. The parameter $\Lambda$ is the spectral width of the reservoir. The atom transition frequency and the cavity are in resonance with $\omega_{\circ}=\omega_{\rm{cavity}}$. When the $\Lambda$ becomes very large or approaches infinity, the reservoir is considered memoryless or Markovian. Then, the cavity is only the main cause of the non-Markovian dynamic in the system. The diagonal (off-diagonal)density matrix elements of the atomic qubit are evolving in time $t$ as a function of $p(t)\propto e^{-\frac{\gamma_{\circ}t}{2}}[cos(\frac{\delta t}{4})+\frac{\gamma_{\circ}}{\delta}sin(\frac{\delta t}{4})]^{2}$ ($\sqrt{p(t)}$), respectively, where $\delta=\sqrt{16\kappa^{2}-\gamma_{\circ}^{2}}$, and $\kappa$ is the coupling constant between the atomic qubit and the cavity and related to the decay rate of the excited state. The parameter $\delta$ can be used to determine whether the system is Markovian ( $4\kappa<\gamma_{\circ}$) or non-Markovian ($4\kappa\geq\gamma_{\circ}$).
The relationship of $\gamma_{\circ}$ and $\kappa$ on the Markovian and non-Markovian processes. The density matrix of the photon state is dependent on the probability function Chen_PRA_14 ; Wang_scr_17 $p(t)\propto e^{-\frac{\gamma_{\circ}t}{2}}[cos(\frac{\delta t}{4})+\frac{\gamma_{\circ}}{\delta}sin(\frac{\delta t}{4})]^{2}$, where $t$ is the propagation time of the photon in the fiber. The parameter $\delta=\sqrt{16\kappa^{2}-\gamma_{\circ}^{2}}$ can be used to determine whether our open quantum system is Markovian or non-Markovian.
For $4\kappa<\gamma_{\circ}$ or $1<\frac{\gamma_{\circ}}{4\kappa}$, $\delta\rightarrow i\delta$, the harmonic function $sin(\cdot)$ and $cos(\cdot)$ becomes hyperbolic function $cosh(\cdot)$ and $sinh(\cdot)$, respectively. Our quantum system is undergoing Markovian dynamics. This occurs when the PMD and the loss parameter are the dominant effects on the photonic qubit. Hence, the probability function $p(t)$ becomes an exponential function. The polarization mode dispersion (PMD) effect that leads to the Markovian process i.e. entanglement sudden death has been experimentally observed Brodsky_OL_2011 and theoretically studied Brodsky_PRL_2011 on the fiber-based entangled photon pair. One can enhance the PMD effect on photon pairs by increasing their differential group delays such as inserting a few polarization-maintaining fiber jumpers in their channels.
For $4\kappa\geq\gamma_{\circ}$, where the PMD effect is less dominant than the loss parameter, our system becomes non-Markovian. For simplicity, we assume $\gamma_{\circ}=4\kappa/n$ where $n>1$, then $\delta=4\kappa\sqrt{1-\frac{1}{n^{2}}}$. The probability function of the qubit is given by,
case 1: $n=\sqrt{2}$
$p_{1}(t)\propto e^{-\frac{\gamma_{\circ}}{2}t}[\rm{cos}(\frac{\kappa_{1}}{\sqrt{2}}t)+\rm{sin}(\frac{\kappa_{1}}{\sqrt{2}}t)]^{2}.$ (8)
case 2: large n
$p_{2}(t)\propto e^{-\frac{\gamma_{\circ}}{2}t}[\rm{cos}(\kappa_{2}t)]^{2}.$ (9)
Then, the harmonic function of $p_{1,2}(t)$ depends only on the PMD effect. However, our analytical approach can provide a new physical insight into the coupling constant ($\kappa$) of the PMD and the loss ($\mu$) effects on the two-photon polarization entangled state when one photon from the pair is stored in a fiber-based quantum buffer.
Amplitude Damping on one photon from the two-photon Werner State. The amplitude damping process on one qubit from two-qubit Werner state $\rho$ is given by,
$\rho^{\prime}=E_{\circ}\rho{E_{\circ}}^{\dagger}+E_{1}\rho{E_{1}}^{\dagger}=\begin{pmatrix}\frac{1+p}{4}&0&0&\frac{p\sqrt{1-\xi}}{2}\\ 0&\frac{(1-p)(1-\xi)}{4}+\frac{(1+p)(\xi)}{4}&0&0\\ 0&0&\frac{1-p}{4}&0\\ \frac{p\sqrt{1-\xi}}{2}&0&0&\frac{(1+p)(1-\xi)}{4}+\frac{(1-p)(\xi)}{4}\end{pmatrix}$
where $E_{\circ}=(I\otimes\Gamma_{\circ})$ and $E_{1}=(I\otimes\Gamma_{1})$. The identity matrix $I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$. The amplitude damping operators are $E_{\circ}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\xi}\end{pmatrix}$ and $E_{1}=\begin{pmatrix}0&\sqrt{\xi}\\ 0&0\end{pmatrix}$, where the $\xi$ is the probability of depolarizing a photon. We measure the density matrix $\rho^{\prime}$ and obtain the density matrix elements $\rho^{\prime}_{11}=\frac{1+p^{\prime}_{11}}{4}$, $\rho^{\prime}_{22}=\frac{1-p^{\prime}_{22}}{4}$, $\rho^{\prime}_{33}=\frac{1-p^{\prime}_{33}}{4}$, $\rho^{\prime}_{44}=\frac{1+p^{\prime}_{44}}{4}$, $\rho^{\prime}_{14}=\frac{p^{\prime}_{14}}{2}$, and $\rho^{\prime}_{41}=\frac{p^{\prime}_{41}}{2}$. From each density matrix element, we write ${p^{\prime}_{nm}}_{(n,m=1,2,3,4)}$ in terms of p of the Werner State $\rho$ such as;
$p^{\prime}_{11}$ $=$ $p^{\prime}_{33}=p$ (10)
$p^{\prime}_{22}$ $=$ $p^{\prime}_{44}=p(1-2\xi)$ (11)
$p^{\prime}_{14}$ $=$ $p^{\prime}_{41}=p\sqrt{1-\xi}\simeq p(1-\xi).$ (12)
where we approximate $p\sqrt{1-\xi}\simeq p(1-\xi)$ for $\xi<0.04(4\%)$. Our experimental data indicates that p drops less than $4\%$. We then take the average of the measured $\frac{p^{\prime}_{11}+p^{\prime}_{22}+p^{\prime}_{33}+p^{\prime}_{44}+p^{\prime}_{14}+p^{\prime}_{41}}{6}=\frac{6p-5p\xi}{6}\sim p(1-\xi)=p(t)$. The $p(t)$ is the non-Markovian probability function. It shows that the p of the Werner state is changing by $(1-\xi)$.
The probability function for $\Delta\phi_{h}\neq\Delta\phi_{v}$ For the situation $\Delta\phi_{h}\neq\Delta\phi_{v}$, Eq.(2) can be rewritten as
$\mathcal{P}_{\pm}^{f}$ $=$ $e^{-2\mu L}[2+2\,\rm{cos(\Delta\phi_{v})}\rm{cos(\Delta\phi_{h})}$ (13)
-- $2\,\rm{sin(\pm\Delta\phi_{h})}\rm{sin(\pm\Delta\phi_{v})}-2\,\rm{cos(\Delta\phi_{h})}\rm{sin(\pm\Delta\phi_{v})}$
$+$ $2\,\rm{cos(\Delta\phi_{h})}\rm{sin(\pm\Delta\phi_{h})}-2\,\rm{cos(\Delta\phi_{v})}\rm{sin(\pm\Delta\phi_{v})}$
$+$ $2\,\rm{cos(\Delta\phi_{v})}\rm{sin(\pm\Delta\phi_{h})}].$
To simplify the above expression, we assume $\Delta\phi_{h(v)}$ is relatively small and substitute $\rm{cos(\Delta\phi)}=1-\frac{\Delta\phi^{2}}{2}$ and $\rm{sin(\Delta\phi)}=\Delta\phi$ to the $\mathcal{P}_{\pm}^{f}$. The first few terms of the $\mathcal{P}_{\pm}^{f}$ are given as;
$\mathcal{P}_{\pm}^{f}\propto e^{-2\mu L}[2+\pm\mathcal{B}L^{\frac{1}{2}}\pm\mathcal{B^{\prime}}L\mp\mathcal{C}L^{\frac{3}{2}}\pm\mathcal{C^{\prime}}L^{2}\pm\mathcal{C}L^{\frac{5}{2}}\pm..].$ (14)
This approximated probability function is depending on the power of the fiber length $L^{\frac{n}{2}}$, where the n is positive integer number.
V Data availability
The data that supports the findings of this study are available from the corresponding author upon reasonable request.
Acknowledgements.